OpenStudy (anonymous):

Find a rational zero of the polynomial function and use it to find all the zeros of the function. f(x) = x^4 + 3x^3 - 5x^2 - 9x - 2

OpenStudy (anonymous):

are u familiar with rational root theoreom?

OpenStudy (anonymous):

yeah I think it is -1,2,-2 but my choices have {1,-2,-2 + sqrt3, -2-sqrt3} this is just an example there is four of them so that confused me

OpenStudy (anonymous):

a. {1,-2,-2 + sqrt3, -2-sqrt3} b. {-1,-2,-2 + sqrt5, -2-sqrt5} c. {-1,3,-2 + sqrt5, -2-sqrt5} d. {-1,2,-2 + sqrt3, -2-sqrt3}

OpenStudy (anonymous):

-1 , 2 is only rational roots

OpenStudy (anonymous):

d is right

OpenStudy (anonymous):

or so after that I have to find zero

OpenStudy (anonymous):

u must factorise after that to find irrational zeros

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

i know that x=-1 is a root so i can factor out (x+1) like this \( x^4 + 3x^3 - 5x^2 - 9x - 2=x^4+x^3+2x^3+2x^2-7x^2-7x-2x-2\\=x^3(x+1)+2x^2(x+1)-7x(x+1)-2(x+1)=(x+1)(x^3+2x^2-7x-2) \)

OpenStudy (anonymous):

x=2 is a root also so u try to factor out x-2 from degree 3 polynomial , finally u will have a quadratic

OpenStudy (anonymous):

so you just place 2 where 1 is right